OpenStudy (anonymous):

How do I know if this solution is extraneous? Please help. sqrt {x+7}-x=1

OpenStudy (anonymous):

\[\sqrt{x+7}-x=1\]

OpenStudy (kc_kennylau):

\[\begin{array}{rcl} \sqrt{x+7}-x&=&1\\ \sqrt{x+7}&=&1+x\\ x+7&=&x^2+2x+1\\ x^2-x-6&=&0\\ (x-3)(x+2)&=&0\\ x-3=0&or&x+2=0\\ x=3&or&x=-2 \end{array}\]

OpenStudy (kc_kennylau):

Substitute both solutions back to the equation and see if the equation is valid

OpenStudy (kc_kennylau):

If the equation is not valid, then that solution is extraneous.

OpenStudy (jack1):

i think... sqrt {x+7}-x = 1 sqrt {x+7} = 1+ x (sqrt {x+7})^2 = ( 1+ x )^2 x+7 = 1 + 2x + x^2 0 = 1 + 2x + x^2 - x - 7 0 = -6 + x + x^2 0 = (x-2)(x+3) so x = 2 or x = -3 pick which is right by substitution

OpenStudy (jack1):

*sigh*... u types whay quicker @kc_kennylau

OpenStudy (jack1):

;D

OpenStudy (kc_kennylau):

@Misskatierae Please give @Jack1 a medal because he answered first, I just typed quicker than him.

OpenStudy (jack1):

nah, fastest finger first @Misskatierae , give it to @kc_kennylau

OpenStudy (kc_kennylau):

@Misskatierae yep medal him you've made the correct decision :)

OpenStudy (jack1):

lol, i think we're spamming u're Q now @Misskatierae , sorry ;D cheers kc

OpenStudy (anonymous):

How about whoever can explain how they got the answer gets a medal? Because I'm still confused haha

OpenStudy (jack1):

u go @kc_kennylau , it'll be quicker ;D

OpenStudy (anonymous):

Haha!! @kc_kennylau can you please help?

OpenStudy (kc_kennylau):

Wait my answer was wrong Jack's answer was correct

OpenStudy (anonymous):

@Jack1 can you please help? lol

OpenStudy (kc_kennylau):

When x=2, L.H.S. =sqrt(2+7)-2 =sqrt9-2 =3-2 =1 R.H.S. =1 \(\therefore\)x=2 is not extraneous

OpenStudy (kc_kennylau):

When x=-3, L.H.S. =sqrt(-3+7)-2 =sqrt4-2 =2-2 =0 R.H.S. =1 \(\therefore\)x=-3 is extraneous.

OpenStudy (kc_kennylau):

@Misskatierae please give @Jack1 a medal since he answered correctly :)

OpenStudy (jack1):

hahaha, most gentlemanly answer string ever ;D

OpenStudy (kc_kennylau):

@Jack1 yep I agree xD

OpenStudy (jack1):

wait, does that all make sense now @Misskatierae ? sidebar: (attached)

OpenStudy (anonymous):

nope @Jack1

OpenStudy (jack1):

ok... which part are you getting lost at, will slow it down and go into detail

OpenStudy (anonymous):

I don't get any of it haha @Jack1 Thank you both so much for helping though!

OpenStudy (kc_kennylau):

Complete solution: \[\begin{array}{rcl} \sqrt{x+7}-x&=&1\\ \sqrt{x+7}&=&x+1\\ x+7&=&x^2+2x+1\\ x^2+x-6&=&0\\ (x+3)(x-2)&=&0\\ x+3=0&\mbox{or}&x-2=0\\ x=-3&\mbox{or}&x=2 \end{array}\]When x=-3, \[L.H.S.\\=\sqrt{-3+7}-(-3)\\=\sqrt4+3\\=7\\R.H.S.\\=1\\\because L.H.S.\ne R.H.S.\\\therefore x=-\mbox{3 is extraneous.}\]When x=2, \[L.H.S.\\=\sqrt{2+7}-2\\=\sqrt9-2\\=3-2\\=1\\R.H.S.\\=1\\\because L.H.S.=R.H.S.\\\therefore x=\mbox{2 is not extraneous.}\]

OpenStudy (anonymous):

Thank you soo much @kc_kennylau!!

OpenStudy (jack1):

nice use of LaTex too dude @kc_kennylau

OpenStudy (kc_kennylau):

lolz thanks @Jack1 :D